1) The sum of three consecutive numbers is 1305. Find the three numbers.
2) The sum of two consecutive natural numbers:

Answer by rfer(16322)

x+x+1+x+2=1305
3x+3=1305
3x=1302
x=434
x+1=435
x+2=436